# On the Robustness of Fully-Spiking Neural Networks in Open-World Scenarios using Forward-Only Learning Algorithms

### v1

@@ 3.3 Interpretation of Latent Vector
Our FF-SCP algorithm presents a method to provide robustness guarantees for AI products. However, this OoD detection method only offers a quantitative measurement indicating the extent to which a sample does not belong to the training dataset. To enhance the functionality of this algorithm, this section introduces a novel interpretability algorithm designed to produce human-understandable explanations of the decisions made by FF-SCP. Given an input sample $x$, this algorithm generates an attribution map highlighting the features that do not align with those expected from the data distribution of a selected class $c$. This method leverages the same latent space approximation used by FF-SCP, employing a large sample set of latent vectors to approximate the emerging topology from the network’s activity (see initialize_latents in Algorithm 1). By exploring this latent space, our algorithm identifies the closest neighbor $x_{f}$ of the sample $x$, ensuring that its latent representations closely resemble those expected from the latent data of class $c$ while retaining key features that characterize the original input $x$. A visual representation of this process is illustrated in Figure 2.
To search the latent space for the desired ID sample $x_{f}$, we employ a decoder network that accurately maps the latent space back to its pre-image space in $D_{\oplus}$. Networks trained using FFA are known to provide accurate projections of the input images, capturing the key distinctive features of this input space, thus allowing for precise reconstruction of the latents [5]. Consequently, we can assert that latent vectors surrounding the latent class manifold will result in plausible reconstructions, not distant from those present in the training dataset. We denote the application of this decoder network as $\texttt{Decode}:\mathbb{L}\rightarrow D_{\oplus}$, which is trained using the input-latent pairs from the positive dataset. It’s important to note that while negative samples belong to the latent space of the network and provide accurate representation probabilities, they do not align with the features expected for the ID data. Therefore, to obtain the features that cause an input sample to deviate from the ID of a specific class, it is important to create a decoder that operates exclusively within the positive distribution.
Our algorithm operates by employing this decoder network to explore the obtained latent space approximation to find an alternative sample $x_{f}$, derived from the input $x$, so that it better fits the input distribution of a user-selected class $c\in\mathcal{C}$. This exploration is achieved by using a parameter latent vector $\ell_{w}$, whose reconstruction results in our desired input vector $x_{f}$. This vector is optimized using a gradient descent algorithm to minimize two key distances: first, the Latent Distance, which measures the displacement between the parameter latent $\ell_{w}$ and the manifold representing the latent cluster $\mathcal{L}_{c,c}$; and second, the Decoder Distance, which measures the distance between the pre-image $x_{f}$ and the original sample $x$. This loss function is encapsulated in the following expression:
@@
where $\alpha$ denotes a hyperparameter used to balance the strength of each distance. Low values of $\alpha$ result in latent values closer to the original image, which are desirable in cases where the overall regions do not account for a large portion of the input image. Conversely, higher values of $\alpha$ result in pre-images of the latent vector being closer to the input distribution of the target class, which is crucial when the distortion of the input data is large.
Once the optimization process finishes and the latent vector $\ell_{w}$ has reached an equilibrium point between the desired targets, the reconstructed sample $x_{f}$ should result in the closest instance from the input space resembling the original image while belonging to the desired class. By using the difference between the original sample $x$ and the optimal instance $x_{f}$, the obtained difference map would highlight the regions that are pushing the sample outside the desired class distribution. For instance, when an ID sample is used as the objective in this algorithm, the resulting attribution map should ideally be close to empty, indicating that the decoded latent vector closely resembles the original image with sufficient precision. However, due to the inherent noise of the decoder, some residual attribution differences may still be present. Conversely, when inputting an OoD sample with some artifacts covering the image, the attribution map obtained should highlight these artifacts with positive values, indicating that they do not belong to the ID class distribution. Meanwhile, if the OoD sample has regions of the image missing, the attribution map should highlight these regions with negative values, indicating that an image belonging to the ID distribution should have those missing parts. This final heuristic has been detailed in Algorithm 2.
1
2
Data: Trained decoder on the original ID data, denoted as $\texttt{Decode}(\cdot)$. Objective sample data $x$ and arbitrary class $c\in\mathcal{C}$. Latent space representative sets $\mathcal{L}$.
Result: Attribution map showing the differences from the sample $x$ to the input distribution of the class $c$.
3
4
5 Function get_attribution($x$, $c$, $\alpha$):
6
7      Initialize the parameter latent vector $\bm{\ell}_{w}$ to the value of $\mathcal{N}(x,c)$.
8      while $\ell_{w}$ has not converged do
9             Compute the Decoder Distance from $\|\texttt{Decode}(\ell_{w})-x\|_{2}$
10             Compute the Latent Distance from $\alpha\cdot d(\ell_{w},\mathcal{L}_{c,c})$
11
12            Obtain the final loss $L_{\text{Attr}}$ as the sum of Decoder Distance and Latent Distance
13
14            Update $\bm{\ell}_{w}$ by applying a gradient descent step on $\nabla L_{\text{Attr}}$ ;
15
16      return $x-\texttt{Decode}(\bm{\ell}_{w})$
17
18
Algorithm 2 Obtain Attribution Map
@@ 4 Experimental Setup
    RQ1: Can our proposed sFFA achieve competitive accuracy compared to the original FFA?
@@
    RQ2: Does the enhanced latent representation of sFFA, utilized by our proposed FF-SCP, improve upon the baseline results of SCP and ODIN?
@@
    RQ3: Does the latent space of models trained with sFFA exhibit sufficiently representative properties to develop interpretability mechanisms?
To provide insights into the experimental process used to address these questions, we have organized this section as follows: Section 4.1 outlines the configuration employed for the different models and the training process used to address RQ1; Section 4.2 describes the experimental procedure followed to evaluate our FF-SCP algorithm in relation to RQ2; Section 4.3 presents the process used to assess the quality of the attribution maps generated by our interpretability algorithm to answer RQ3; and finally, Section 4.4 lists the datasets and data processing employed in this work. All the code for the experiments has been uploaded to https://github.com/AnonymousSquirrel316/FFA_OOD.
4.1 Architecture and Training Process
To conduct an unbiased experimental analysis across the different network architectures, minimizing potential biases between analog and spiking networks, all models were designed to achieve equivalent behaviors. Specifically, each model consisted of two densely connected feed-forward layers, each with 1400 neurons. The analog networks employed ReLU activation functions, while the spiking networks relied on LIF neural models as a biological counterpart. To implement these spiking dynamics, we utilized the LIF class from the snnTorch [32] library, with the following hyperparameters: a fast sigmoid as the surrogate function, 20 timesteps per forward pass; a threshold of $0.4$ for the first layer and $0.3$ for the subsequent layer; and $\beta_{n}$ values of $0.4$ for the first layer and $0.2$ for the subsequent layer. To prevent spiking neurons from entering over-excited states, an input downscaling technique was applied. This technique involved scaling down the input to each neuron by a factor of $0.85^{N_{\text{Spikes}}}$ at each timestep, where $N_{\text{Spikes}}$ denotes the total number of output spikes from the respective neuron. This approach effectively reduces the input activities of highly active neurons over time, thereby moderating their overall activity throughout the timesteps.
The training process for all models followed the supervised methodology proposed by Hinton [3], wherein positive and negative samples are generated by embedding an encoded version of the label into the image. In addition to Hinton’s proposed negative generation process, we also employed a greedy negative generation method. In this approach, the negative label is chosen as the one which maximizes the goodness score, aiming at increasing the distance between the goodness scores of positive and negative labels. Each method was independently used to train each network, and the one resulting in higher accuracy was selected for further experimentation to ensure optimal predictive capabilities for subsequent tasks, including OoD detection and our attribution method. Additionally, we adopted the label initialization and embedding techniques proposed by Lee et al. [17], where embedding vectors of class labels are set as random binary vectors appended at the end of the data. Each binary vector, of dimension 100, was initialized using a Bernoulli distribution with a probability of 0.1 for each element of the vector.
All networks were trained for 10 epochs with a batch size of 512, using the hyperparameterized Binary Cross Entropy loss function outlined in Subsection 3.1. The spiking networks were trained using the two goodness functions developed in Section 3.1: the bounded goodness, $G_{0}(\cdot)$, and the unbounded sFFA, $G_{\infty}(\cdot)$. For networks using the unbounded goodness function, the parameters were set as $\theta^{+}=6$ and $\theta^{-}=2$. Networks employing the bounded goodness function used thresholds of $\theta^{+}=0.3$ and $\theta^{-}=0.1$, along with $\alpha=\beta=5$ to optimize the scaling of the probability function. We employed the ADAM optimizer with a learning rate of 0.001 for ANNs and 0.002 for SNNs. These values were determined through a process of early experimentation and manual tuning.
4.2 Out-of-Distribution Evaluation Setup
As this paper builds upon the groundwork of the original SCP algorithm introduced in [24], and given the absence of subsequent spiking-specific OoD detection methods in the literature, the benchmarks used for this work are extracted from the aforementioned study. Specifically, the SCP and ODIN algorithms stand as the top-performing approaches, serving as the designated benchmarks for comparative analysis in this work. Furthermore, to demonstrate the competitiveness of our algorithm in the spiking domain compared to its analog counterpart, both types of networks have been tested.
The models used in these experiments are the best-performing ones from the training experiments of RQ1. We explored the use of Euclidean, Manhattan, and Cosine distances for the distance function $d(\cdot,\cdot)$, and chose the Manhattan distance due to its superior performance. We employed 1024 samples to create the latent space, following the process detailed in the initialize_latents function of Algorithm 1. For the filtering step, we removed 20% of the highest and lowest latent vectors from their respective latent sets. The hyperparameters $\gamma$, $\beta$, and $z_{0}$ were selected through an early hyperparameter selection process for each training and goodness. Their search space was defined as follows: $z_{0}$ was chosen from the set $\{0.85,1,1.15,1.4\}$; $\gamma$ was either $10^{4}$ or $10^{6}$; and $\beta$ was set to either 1 or 2, representing Manhattan or Euclidean aggregation of distances, respectively.
To objectively evaluate the model’s OoD detection performance across the experiments, three key evaluation metrics have been measured: AUROC, AUPR, and FPR95. These metrics are widely employed in the OoD detection literature, providing valuable insights into the model’s ability to distinguish between in-distribution and out-of-distribution samples. Their definitions are as follows:
  • $\circ$
    AUROC, also known as the Area Under the Receiver Operating Characteristic Curve, is a threshold-independent metric that, when given a set of OoD scores for the data, quantifies the extent of overlap between OoD and ID samples. The AUROC metric ranges from 00 to $1$, where $0.5$ indicates a random classifier’s performance, and $1$ signifies a perfect classifier that can perfectly separate OoD and ID samples.
  • $\circ$
    AUPR, also known as the Area Under the Precision-Recall Curve, considers the precision and recall trade-off in classifying OoD samples. AUPR values closer to $1$ indicate better performance, with high precision and recall in distinguishing OoD samples.
  • $\circ$
    FPR95, also known as the False Positive Rate at a 95% True Positive Rate, provides insights into the model’s false positive rate when the true positive rate is set at 95%. This metric quantifies the model’s ability to maintain a low false positive rate while correctly identifying 95% of OoD samples. Lower FPR95 values are indicative of better OoD detection performance.
@@ 4.3 Evaluation of the Quality of the Attribution Method
Figure 3: MNIST and KMNIST dataset samples with applied obstructions for attribution experiments. Each row begins with an in-distribution sample from the respective dataset, followed by its obstructed versions.
@@ 4.4 In-Distribution and Out-of-Distribution Datasets
To evaluate the training accuracy of our adaptation of FFA to spiking networks, we employed five well-known datasets: MNIST [33], consisting on 60,000 images of 10 classes of handwritten digits; Fashion MNIST [34], composed of 60,000 images representing 10 classes of clothing items; Kuzushiji-MNIST (KMNIST) [35], consisting of 60,000 images of Japanese characters divided into 10 different classes; and EMNIST (Letters) [36], composed of 145,000 images of the 26 lowercase letters of the Latin alphabet. A uniform normalization pipeline was applied across all datasets. Each of these spike trains was set to last for 20 timesteps.
Since OoD experiments require a trained model and its corresponding training dataset as the ID baseline, we utilize all the models previously trained in RQ1 for experimentation in RQ2. In these experiments, the OoD dataset is constructed by combining all other training datasets on which the models were not trained. To expand the scope of our analysis, we introduce two additional OoD datasets: Not-MNIST (small) [37], which consists of 19,000 font glyphs representing letters A to J, and Omniglot [38], containing 1,623 handwritten images spanning 50 different alphabets. These datasets underwent the same data processing pipeline used in the training experiments.
4.5 Limitations
This work proposes FFA as a candidate alternative to BP, highlighting its functional latent space and representational capabilities, which offer substantial advantages for developing techniques to enhance model robustness for industry-ready software. However, due to the early stage of research into this algorithm, we observed a clear lack of alternative architectural formulations (e.g., convolutional or transformer layers). To date, only one work has presented a possible design for a convolutional layer in FFA by relying on pattern-like embeddable patterns [39]. While this method demonstrates accurate precision on small datasets, no proofs have been given of this algoritmh scaling to more complex datasets, such as CIFAR-10, where images contain non-empty backgrounds. Consequently, current models can only rely on densely connected layers, which have multiple limitations when learning from highly complex datasets. As a result of this limitation, the latent representations of these models exhibit a clear upper boundary on representational accuracy. These inaccuracies in understanding the data can lead to suboptimal cluster distribution within the latent space, preventing effective leveraging of cluster distances for accurate OoD detection or interpretability mechanisms.
However, it is important to note that the representational properties of FFA arise from its learning dynamics, which are independent of the specific architectural choice of weights. Therefore, this paper provides an initial perspective on the practical utility of FFA in its current form, based on densely connected layers. Nonetheless, as further elaborated in Section 6, one of our primary goals for future research is to develop alternative formulations of FFA suitable for convolutional architectures, thereby extending our findings to more complex datasets.
@@ 5 Results and Discussion
This section presents the results from various experiments, discussing the findings and highlighting observed limitations. Initially, Subsection 5.1 showcases the accuracy results of our proposed spiking adaptations of FFA. Subsection 5.2 presents the outcomes of the OoD detection task using the FF-SCP algorithm, comparing it against the SCP and ODIN algorithms. Finally, Subsection 5.3 discusses the results obtained from the attribution method proposed in this paper.
5.1 RQ1: Training Results of sFFA
The accuracies obtained from the training experiments using the different goodness functions are presented in Table 1. These results demonstrate the comparable performance between the analog and the spiking implementations of FFA, where SNNs only exhibit a slight decrease in accuracy, likely due to the increased noise within latent vectors caused by the stochastic input preparation process. Among the two proposed goodness functions, the models employing the unbounded version achieved higher accuracy than their bounded counterparts, with an accuracy of $86.16\%$ compared to $81.20\%$.
Nevertheless, while the average accuracy shows a noticeable drop, this loss in accuracy is primarily related to the significant difference in accuracy on the MNIST dataset, where the bounded version reduces accuracy by $14.61$ points. This abnormal relationship in the drop in accuracy across different datasets suggests that the bounded and unbounded models process the data differently, leading to distinct classification heuristics and representation properties. These results align with the observation that networks using bounded goodness functions do not experimentally create sparse latent representations, which can result in reduced feature specificity in the neurons (see Appendix B).
Table 1: Accuracy of networks trained with the Forward-Forward Algorithm. The accuracies of models with a similar architecrure (i.e., densely connected layers) and number of parameters have been included for comparison.
Dataset BP          Analog FFA          Analog Spiking FFA $G_{0}(\cdot)$ Spiking FFA $G_{\infty}(\cdot)$
MNIST 98.25% 97.77% 93.27% 93.01%
FMNIST 89.47% 85.75% 83.69% 85.68%
KMNIST 92.11% 91.86% 71.30% 85.91%
EMNIST 92.19% 87.25% 76.54% 80.07%
Average 93.00% 90.65% 81.20% 86.16%

### v2

@@ 3.3 Interpretation of Latent Vector
This section introduces a novel interpretability algorithm aimed at producing human-understandable explanations for the decisions made by FF-OoD. Given an input sample $x$, this algorithm generates an attribution map highlighting the features that do not align with those expected from the data distribution of a selected class $c$. To do so, it relies on the same latent space discrete approximation used by FF-OoD, which accurately approximates the emerging topology from networks trained using FFA (see INITIALIZE_LATENTS in Algorithm 1). By exploring this latent space, our algorithm identifies the closest neighbor $x_{f}$ of the sample $x$, ensuring that its latent representations closely resemble those expected from the latent data of class $c$ while retaining key features that characterize the original input $x$. A visual representation of this process is illustrated in Figure 2.
The search of the latent space for the desired ID sample $x_{f}$ is achieved by employing a decoder network that maps the latent space back to its pre-image space in $D_{\oplus}$. Given the highly representational latent space of networks trained under FFA (Ororbia and Mali, 2023), we can assert that latent vectors surrounding the latent class manifold will result in plausible reconstructions, not distant from those present in the training dataset. We denote the application of this decoder network as $\mathcal{N}^{-1}:\mathbb{L}\rightarrow D_{\oplus}$, which is trained using the input-latent pairs from the positive dataset. It’s important to note that while negative samples belong to the latent space of the network, their features do not align with those expected from ID data. Therefore, to avoid deviating from ID classes, we operate solely within $D_{\oplus}$.
This decoder network allows the algorithm to explore the latent space to find a distinct $x_{f}$, derived from the input $x$, that better aligns with the input distribution of a user-selected class $c\in\mathcal{C}$. This exploration follows an optimization process in which a parameter latent vector $\ell_{w}$ is optimized using gradient descent to minimize two key distances: a Latent Distance, which measures the displacement between the parameter latent $\ell_{w}$ and the manifold representing the latent cluster $\mathcal{L}_{c,c}$; and a Decoder Distance, which measures the distance between the pre-image $x_{f}$ and the original sample $x$. This loss function is encapsulated in the following expression:
@@
where $\alpha$ denotes a hyperparameter used to balance the strength of each distance. Low values of $\alpha$ result in latent values closer to the original image, which are desirable in cases where the sample is close to the desired distribution. Conversely, higher values of $\alpha$ result in pre-images of the latent vector being closer to the input distribution of the target class, which is crucial when the distortion of the input data is large.
Once the optimization process is finished and the latent vector $\ell_{w}$ has reached an equilibrium point between the desired targets, the attribution map is obtained by taking the difference between the original sample $x$ and the optimal instance $x_{f}$. This final heuristic has been detailed in Algorithm 2.
Algorithm 2 Attribution Map Generation
1:  GET_ATTRIBUTION($x$, $c$, $\alpha$)
2:        Initialize the parameter latent vector $\bm{\ell}_{w}$ to $\mathcal{N}(x,c)$.
3:        while $\ell_{w}$ has not converged:
4:            Compute the distance $\|\mathcal{N}^{-1}(\ell_{w})-x\|_{2}$
5:            Compute the distance $\alpha\cdot\widehat{d}(\ell_{w},\mathcal{L}_{c,c})$
6:            Obtain $L_{\text{Attr}}$ as the sum of both distances
7:            Update $\bm{\ell}_{w}$ with a gradient descent step on $\nabla L_{\text{Attr}}$
8:        return $x-\mathcal{N}^{-1}(\bm{\ell}_{w})$
@@ 4 Experimental Setup
  • 1.
    RQ1: Does the surrogate gradient result in competitive accuracy when training spiking FFA?
  • 2.
    RQ2: Does the enhanced latent representation of spiking FFA, utilized by our proposed FF-OoD, improve upon the baseline results of SCP and ODIN?
  • 3.
    RQ3: Does the latent space of models trained with spiking FFA exhibit sufficiently representative properties to develop interpretability mechanisms?
4.1 Architecture and Training Process
To conduct an unbiased experimental analysis across different network architectures and minimize potential biases between analog and spiking networks, all models were designed to exhibit equivalent behaviors. The first set of experiments was conducted on models consisting of two densely connected feedforward layers, each with 1400 neurons. Given the lack of competitive convolutional network implementations using FFA, the second set of experiments utilized a ResNet18 backbone to extract initial features (He et al., 2016), which were then passed to a two-layer fully-spiking network with $512$ neurons per layer. Analog models employed ReLU activation functions, while spike-based networks used LIF neurons as their biological counterpart. To implement these spiking dynamics, we used the LIF class from the snnTorch (Eshraghian et al., 2023) library, with the following hyperparameters: a fast sigmoid as the surrogate function, $20$ timesteps per forward pass; a threshold of $0.4$ for the first layer and $0.3$ for the subsequent layer; and $\beta_{n}$ values of $0.4$ for the first layer and $0.2$ for the subsequent layer. To prevent spiking neurons from entering over-excitepd states, an input downscaling technique was applied, where each neuron’s input was scaled by a factor of $0.85^{N_{\text{Spikes}}}$ at each timestep, with $N_{\text{Spikes}}$ representing the total number of output spikes from that neuron at that moment.
The training process for all models followed the supervised methodology proposed by Hinton (2022), wherein positive and negative samples are generated by embedding an encoded version of the label into the data. In addition to Hinton’s proposed negative generation process, we also employed a greedy negative generation method, where the training negative label is chosen to the one showcasing the maximal goodness score. Each method was independently used to train each network, and the one resulting in higher accuracy was selected for further experimentation to ensure optimal predictive capabilities for subsequent tasks, including OoD detection and our attribution method. Additionally, we adopted the label initialization and embedding techniques proposed by Lee and Song (2023), where embedding vectors of class labels are set as random binary vectors appended at the end of the data. Each binary vector, of dimension 100, was initialized using a Bernoulli distribution with a probability of 0.1 for each element of the vector.
Networks without a backbone were trained for 10 epochs with a batch size of 512, using the hyperparameterized Binary Cross Entropy loss function outlined in Equation 6. Spiking networks were trained with the goodness functions $G_{0}(\cdot)$ and $G_{\infty}(\cdot)$. For networks utilizing the unbounded goodness function, the parameters were set to $\theta^{+}=6$ and $\theta^{-}=2$. Networks employing the bounded goodness function used thresholds of $\theta^{+}=0.3$ and $\theta^{-}=0.1$, along with $\alpha=\beta=5$ to optimize the probability function’s scaling. Training was performed using the ADAM optimizer with a learning rate of 0.001 for ANNs and 0.002 for SNNs. In contrast, networks using a ResNet backbone were trained for 100 epochs. For unbounded models, the parameters were set to $\theta=12$ with $\alpha$ and $\beta$ both equal to 5. For bounded models, $\theta$ values were set to 0.1 for CIFAR-10 and 0.5 for SVHN, with $\alpha$ and $\beta$ set to 10 for CIFAR-10 and 4 for SVHN. These values were determined through early experimentation and manual tuning. All the code has been uploaded into https://github.com/AnonymousSquirrel316/FFA_OOD.
4.2 Out-of-Distribution Evaluation Setup
As this paper builds upon the groundwork of the SCP algorithm introduced by Martinez-Seras et al. (2023), and given the absence of subsequent spiking-specific OoD detection methods in the literature, the benchmarks used for this work are extracted from the aforementioned study. Specifically, the SCP and ODIN algorithms stand as the top-performing approaches, serving as the designated benchmarks for comparative analysis in this work. Furthermore, to demonstrate the competitiveness of our algorithm in the spiking domain compared to its analog counterpart, both types of networks have been tested.
The models used in these experiments were the best-performing ones from the training phase of RQ1. During early experimentation, we explored Euclidean, Manhattan, and Cosine distances for the distance function $d(\cdot,\cdot)$ and selected Manhattan distance for non-backbone networks and Euclidean distance for networks using a feature extractor, as they yielded superior performance. To construct the latent space, we used 1024 samples for networks without a backbone and 5000 samples for networks with a backbone, following the procedure outlined in the initialize_latents function of Algorithm 1. For the filtering step, we removed 20% of the highest and lowest latent vectors from their respective latent sets. The hyperparameters $\gamma$, $\beta$, and $z_{0}$ were selected through an early hyperparameter selection process for each training and goodness. Their search space was defined as follows: $z_{0}$ was chosen from the set $\{0.85,1,1.15,1.4\}$; $\gamma$ was either $10^{4}$ or $10^{6}$; and $\beta$ was set to either 1 or 2.
To objectively evaluate the model’s OoD detection performance across the experiments, three key evaluation metrics have been measured: AUROC, AUPR, and FPR95. These metrics are widely employed in the OoD detection literature, providing valuable insights into the model’s ability to distinguish between in-distribution and out-of-distribution samples. They are defined as:
@@ 4.2 Out-of-Distribution Evaluation Setup
    AUROC (Area Under the Receiver Operating Characteristic curve): Quantifies the model’s ability to discriminate between in-distribution and out-of-distribution samples. A higher AUROC indicates better separation, with 1.0 representing perfect discrimination and 0.5 indicating random guessing.
@@
    AUPR (Area Under the Precision-Recall curve): Evaluates the trade-off between precision and recall for detecting OoD samples. It is particularly useful when dealing with imbalanced datasets, as it emphasizes the model’s ability to correctly identify OoD instances.
@@
    FPR95 (False Positive Rate at 95% True Positive Rate): Measures the false positive rate when the true positive rate is fixed at 95%. Lower FPR95 scores indicate higher OoD detection performance.
@@ 4.3 Evaluation of the Quality of the Attribution Method
Figure 3: MNIST and KMNIST ID samples followed by OoD version obtained by applying one obstruction method.
@@ 4.4 In-Distribution and Out-of-Distribution Datasets
To evaluate the OoD Detection capacity of spiking networks, we trained the networks in 5 ID datasets for the feedforward (MNIST (LeCun and Cortes, 2010), Fashion MNIST (FMNIST) (Xiao et al., 2017), Kuzushiji-MNIST (KMNIST) (Clanuwat et al., 2018), and EMNIST (only the Letters subset) (Cohen et al., 2017)) and 2 for the convolutional alternative (CIFAR10 (Krizhevsky et al., 2009) and SVHN (Netzer et al., 2011)). In these experiments, the OoD dataset is constructed by combining the datasets from each category excluding the one used for the training. To expand the scope of our analysis, we introduce two additional OoD datasets for the densely connected model (Not-MNIST (small) (Bulatov, 2011) and Omniglot (Lake et al., 2015)) and six for the convolutional (Food101 (Bossard et al., 2014), Flowers102 (Nilsback and Zisserman, 2008), Caltech101 (Fei-Fei et al., 2006), GTSRB (Stallkamp et al., 2011), EuroSAT (Helber et al., 2019) and DTD (Cimpoi et al., 2014)). For the densely connected model, a uniform normalization pipeline was applied across all datasets and all images were transformed into 20 timesteps of uniform activity. Similarly, the outputs of the ResNet18 backbone followed the same transformation before going into networks trained using FFA.
@@ 5 Results and Discussion
This section presents the results from various experiments, discussing the findings and highlighting observed limitations. Initially, Section 5.1 showcases the accuracy results of our proposed spiking adaptations of FFA. Section 5.2 presents the outcomes of the OoD detection task using the FF-OoD algorithm, comparing it against the SCP and ODIN algorithms. Subsequently, Section 5.3 discusses the results obtained from the attribution method proposed in this paper. Finally, Section 5.4 discusses the limitations observed during experimentation.
5.1 RQ1: Training Results of spiking FFA
The accuracies obtained from the training experiments using the different goodness functions are presented in Table 1. These results demonstrate the comparable performance between the analog and the spiking implementations of FFA, where SNNs only exhibit a slight decrease in accuracy, likely due to the increased noise within latent vectors caused by the stochastic input preparation process and the information loss due to the discretization of the latent states. Between the two proposed goodness functions, the models employing the unbounded version achieved higher accuracy than their bounded counterparts, probably as a consequence of the probability function not needing additional scaling parameters to reach the extreme values.
Nevertheless, in non-convolutional networks, while the average accuracy shows a noticeable drop, this loss in accuracy is primarily related to the significant difference in accuracy on the KMNIST dataset, where the bounded version reduces accuracy by $14.61$ points. This abnormal relationship in the drop in accuracy across different datasets suggests that the bounded and unbounded models process the data differently, leading to distinct classification heuristics and representation properties. These results align with the observation that networks using bounded goodness functions do not experimentally create sparse latent representations, which can result in reduced feature specificity in the neurons (see C).
When examining networks trained using the output of a feature extractor, the results show that only unbounded goodness remains competitive. The highly variable activity distribution in the feature extractor’s latent space leads to goodness scores that fail to adequately cover the domain of the probability function. This results in suboptimal gradient updates, which eventually converge to less accurate parameters. As a result, these experiments appear to have an accuracy decrease of approximately a 10%.
Table 1: Accuracy of networks trained with the Forward-Forward Algorithm (FFA) using an analog and two spiking implementations. The spiking versions employ the $G_{0}(\cdot)$ and $G_{\infty}(\cdot)$ goodness functions respectively.
Network Dataset FFA           Analog FFA           Spiking $G_{0}(\cdot)$ FFA           Spiking $G_{\infty}(\cdot)$
Dense MNIST 97.77% 93.27% 93.01%
FMNIST 85.75% 83.69% 85.68%
KMNIST 91.86% 71.30% 85.91%
EMNIST 87.25% 76.54% 80.07%
Image CIFAR10 69.86% 61.7% 69.60%
SVHN 90.25% 81.78% 90.57%